OpenStudy (anonymous):

Please help, will medal (and fan depending on how helpful you were ) ^.^ x2 -1 _____ x2 – 9 _____ x2 + 3x – 4 _____________ x2 + 8x + 15

OpenStudy (anonymous):

Do you just need to factor this or what?

Nnesha (nnesha):

for the top part "difference of square" use this method and for bottom part factor that first

Nnesha (nnesha):

difference of square \[\huge\rm a^2 - b^2 = (a+b)(a-b)\]

Nnesha (nnesha):

so what is \[\huge\rm (x^2 -1) = ????\]

OpenStudy (anonymous):

wow I'm already lost

OpenStudy (anonymous):

@Tuggaro I'm not actually sure, it was just this big stack, I think they just want me to simplify it as much as I can

Nnesha (nnesha):

 this is ur question